# Fractional Chern insulator candidate in twisted bilayer checkboard lattice

### v1

@@ Abstract
We investigate a fractional Chern insulator (FCI) candidate arising from Moiré bands with higher Chern number $C=2$ on a magic angle twisted bilayer checkboard lattice (MATBCB). There are two nearly flat low lying bands in the single particle energy spectrum under the first magic angle $\phi\approx 1.608^{\circ}$ and chiral limit. We find MATBCB hosts a nearly uniform Berry curvature distribution and exhibits tiny violation of quantum geometric trace condition in the first moiré Brillourin Zone (mBZ), indicating that there is a nearly ideal quantum geometry in MATBCB in single particle level. Turning on projected Coulomb interactions, we perform exact diagonalization and find a ten-fold ground state quasi-degeneracy in many body energy spectrum with filling fraction $\nu=1/5$. The ten-fold quasi-degenrate ground states further show spectra flow under flux pumping. By diagnosing the particle entanglement spectrum (PES) of the ground states, we obtain a clear PES gap and quasi-hole state counting consistent with Halperin spin singlet generalized Pauli principle, suggesting that a fractional Chern insulator is realized in this system.
June 13, 2024
@@ I Introduction
These experiments have raised a great interest among theorists. Several theories have been proposed to interpret the FQAHE (zero magnetic field FCI) in $\rm{MoTe_{2}}$ [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and R5G [24, 25, 26, 27, 28, 29, 30]. Besides gapped phases, gapless phases beyond FQAHE, e.g., the composite fermi liquid phase, have also been found [31]. However, the experiments mentioned above all seem to focus on fractional filling $C=1$ Chern band which has well known lowest Landau level (LLL) correspondence. When the topological flat band host a higher Chern number, its potential FCI phase will not have a naive LLL correspondence [32, 33, 34, 35]. What is more attracting, it is possible to realize non-abelian FCI phase [36, 32, 37, 38, 39, 40, 41, 20, 42, 43, 44] in fractionally filled flat band with higher Chern number, which could be a building block for fault-tolerant quantum computation.
Motivated by above progress, it is valuable to search for higher Chern number FCI phase in real material or artificial system. In this manuscript, we theoretically study a $C=2$ FCI candidate in magic angle twisted bilayer checkboard (MATBCB) lattice model [45] with fractional filling $\nu=1/5$ of the lowest single particle conductive band. By using exact diagonalization (ED) in momentum space, we find there is a ten-fold ground state quasi-degeneracy in the many-body energy spectrum and the spectrum flow under adiabatic flux inserting. These ten-fold low energy states are the topological ground state candidates for putative $C=2$ FCI phase. From the perspective of many-body entanglement, we study the particle entanglement spectrum (PES) of these ten-fold ground states density matrix. We find that the quasi-hole counting under PES gap obeys the Halperin spin singlet generalized Pauli principle (GPP), which is a smoking gun evidence for the existence of $C=2$ FCI phase. This higher Chern number FCI phase may be realized in artificial quantum simulator such as Rydberg atom array [46], twisted optical lattice [47, 48, 49, 50], circuit QED system [51]. Up to now, we have noticed that there are several works have studied TBCB flat band [45], or similar system like twisted square lattice [52, 53, 54, 55], twisted FeSe [56], strained moiré lattice with quadratic dispersion [57], and square lattice moiré heterostructure [58].
@@ II Single particle moiré band structure of MATBCB
In this section we solve the Moiré Hamiltonian (free fermion hopping in momentum space) of MATBCB and show the flat band structure, as well as the non-trivial Chern number. The Moiré Hamiltonian for a given momentum $H(k_{x},k_{y})$ expressed in reciprocal lattice basis is shown in Eq. (1):
@@
Up to now we have a preliminary picture of MATBCB in single particle level including single particle band and Wilson loop winding. In the next section, we will use quantum geometry indicators to inspect if MATBCB is flavor to host FCI phase more carefully.
@@ III Quantum geometry indicators for fractional Chern insulator
Figure 2: (a) The Wilson loop winding of the lowest two flat bands near half filling. Since the single particle band is gapless at X, the Wilson loop is a $2\times 2$ matrix. The windings are corresponding to the two eigenvalues of the Wilson loop. One can read out the Chern number $C=\pm 2$. From the variation speed of the winding, one can believe that the Berry curvature distribution is uniform enough to host a FCI phase. (b) The scheme of the k point discretization in the first mBZ. To avoid the boundary discontinuity one should take one more round of the k points into account. For example, for $3\times 3$ k point discretization, one has to consider $4\times 4$ k points as shown. (c) The Berry curvature given by QGT. The k point density in the first mBZ is $74\times 74$. (d) The trace of quantum metric in the first mBZ using the same k point density. It almost has the same distribution with Berry curvature, which implies the single particle flat band Bloch function is close to a holomorphic function about $k=k_{x}+ik_{y}$ and ideal trace condition violation. One can also inspect the Berry curvature given by Wilson loop Eq. 11 as benchmark, which is almost the same as (c). The range of momentum $k_{x},k_{y}$ are the same $[0,2\pi]$ for (c) and (d).
@@
In fact, strictly speaking, one should keep in mind that the quantum geometry indicators mentioned above are only directly applicable to the $C=1$ case which has LLL correspondence. Fortunately, several recent works have shown that ideal higher Chern band can be decomposed into a set of $C=1$ ideal Chern bands [65, 66] under certain condition. Specially, when the Berry curvature of $C>1$ Chern band satisfies $F(\textbf{k})=F(\textbf{k}+\textbf{Q})$ with Q being some translation symmetry broken vector, the set of decomposed $C=1$ ideal Chern bands can form orthogonal basis. That means these $C=1$ bands can be independently filled. So the $C=1$ quantum geometry indicators can be used as usual. In our MATBCB case, $\textbf{Q}=(G_{1}/2,G_{2}/2)$. One can see this point later in detail. A more general FCI indicator may be the so- called vortexability [67, 68], which we leave for future work.
@@
In practice, one can use the following simplified form of QGT Eq. (10). To avoid the possible boundary discontinuity [33], one has to compute one another round of k points beyond the first mBZ as shown in Fig. 2. They act as the similar effect like embedding matrix [45]. The profile of Berry curvature is similar to the result given by Ref. [69], which indicates topological heavy fermion is also a potential competing phase.
$\begin{split}\eta^{ab}_{\mu\nu}(\textbf{k})\approx\frac{\langle u^{a}_{\textbf{k}+dk_{\mu}}|u^{b}_{\textbf{k}+dk_{\nu}}\rangle-\sum_{n}\langle u^{a}_{\textbf{k}+dk_{\mu}}|u^{n}_{\textbf{k}}\rangle\langle u^{n}_{\textbf{k}}|u^{b}_{\textbf{k}+dk_{\nu}}\rangle}{dk_{\mu}dk_{\nu}}.\end{split}$ (10)
One can inspect the Berry curvature from QGT and Wilson loop. Compared with the trace of quantum metric, one can see how large the violation of trace condition is. The Berry curvature and Chern number from Wilson loop formalism are shown in Eq. (11). For the mBZ (torus $T^{2}$), there is no need to fix the gauge [70].
$\begin{split}&[U_{r}(\textbf{k})]_{ab}=\langle u_{a}(\textbf{k})|u_{b}(\textbf{k}+d\textbf{k}_{r})\rangle,r=x,y.\\ &W(\textbf{k})=[U_{y}(\textbf{k})]^{-1}[U_{x}(\textbf{k}+dk_{y})]^{-1}U_{y}(\textbf{k}+dk_{x})U_{x}(\textbf{k})\\ &=\exp(iF(\textbf{k})dk_{x}dk_{y}),\\ &W=USV^{\dagger},\quad W\to UV^{\dagger},\\ &C=\frac{1}{2\pi}\int_{mBZ}d^{2}kTr(F)=\frac{1}{2\pi i}\int_{mBZ}Tr(\ln(W(\textbf{k})))\\ &=\frac{1}{2\pi i}\int_{mBZ}\ln(\det(W(\textbf{k})))\end{split}$ (11)
From Fig. 2, one can see that Berry curvature of $C=+2$ band (Wilson eigenvalue) from QGT and the trace of quantum metric are consistent, which means the ideal trace condition violation condition is well satisfied. More importantly, one can see that the Berry curvature host the symmetry $F(\textbf{k})=F(\textbf{k}+\textbf{Q})$, $\textbf{Q}=(G_{1}/2,G_{2}/2)$. That means the $C=2$ band can be orthogonally decomposed into two $C=1$ bands. Therefore the original quantum geometry indicator for $C=1$ case is still valid in our $C=2$ MATBCB system. For MATBCB, we summarize the quantum geometry indicators as follow:
@@
    The standard deviation of Berry curvature is $\sigma[\eta]=6.013$.
@@
    The trace condition deviation is $T[\eta]=0.0149$.
Comapared with MATBG [13, 64], MATBCB has more ideal bandwidth and trace condition violation but slightly worse Berry curvature deviation. We note that some recent works have proposed the concept of higher Berry curvature [71, 72, 73], which may develop to many-body version quantum geometry. How to apply it to moiré system will be an interesting problem.
Figure 3: The variation of trace condition violation $T[\eta]$ over chiral ratio $\kappa$. One can read out that $T[\eta]$ is monotonically increasing with $\kappa$, which indicates that non-zero chiral ratio can induce the instability of putative FCI phase.
@@ IV Many-body energy spectrum and spectrum flow for FCI candidate
Figure 5: (a) The many body energy spectrum for $\nu=1/5$ MATBCB under chiral limit. The discretization of the first mBZ is $4\times 5$. The number of mBZ is $7\times 7$ which is the size of quadratic nodes array. We will use the same parameters in the following simulation unless specified explaination. (b) The zoomed in of low energy window in (a). One can identify the ground state manifold with GSD=10 at k sector $K=5,6\cdots 9$ and $K=15,16\cdots 19$. They are approximately related by vector $G_{1}/2$. The gap between ground state manifold and the higher excitation states is about 12 meV.
@@
Figure 6: (a) The spectrum flow under adiabatic flux insertion $k_{y}\to k_{y}+\Phi/N_{y}$. (b) The zoomed in of low energy window in (a). One can examine that the states related by $G_{1}/2$ will keep almost two-fold degeneracy during flux inserting. The ten-fold quasi-degenerate ground states braid with each other as flux varying, while do not mix with higher energy states. These 10 states can be considered as the FCI topological ground state candidates. The gap between ground state manifold and higher energy states is about $5-12{\rm meV}$, While the width of ground state manifold is approximately $1-2{\rm meV}$.
@@ V The particle entanglement spectrum for MATBCB and quasi-hole counting
Figure 7: The PES of ten-fold ground state density matrix for $\kappa=w_{AA}/w_{AB}=0$. The PES is labelled by the total momentum of subsystem A $K^{A}$. The spectrum almost gather at $\xi<16$ regime, while other data with higher values can be considered as noise. There is an apparent PES gap at $\xi\approx 6.5$. One can extract the PES quasi-hole counting at each $K^{A}$ sector under the PES gap. For $K^{A}=0,1\cdots 4$, there are 4 states per sector. While for $K^{A}=5,6\cdots 19$, there are 5 states per sector. The total number of quasi-hole is 95 (for $4\times 5$ momentum space size). The counting obeys the $(1,5)_{C=2}$ GPP based on Halperin spin singlet ansatz as argued in main text.
@@
From the above analysis, now one can come back and inspect the PES of $\nu=1/5$ MATBCB. In Fig. 7, we show the PES of the density matrix consists of 10 fold topological ground state candidates. There is an apparent PES gap at about $\xi=6.5$ as shown by the blue dashed line. One can check that there are 4 quasi-hole states under PES gap in each $K^{A}=0,1\cdots 4$ sector while 5 low quasi-hole states in each $K^{A}=5,6\cdots 19$ sector. The total number of quasi-hole state is 95 for $N_{x}=4,N_{y}=5,N_{A}=2$. Such quasi-hole counting is coincident with $(1,5)_{C=2}$ GPP. One can interpret this result as follows. In our case, $N_{x}=4,N_{y}=5,N_{e}=4,N_{A}=2$. One can enumerate the Halperin spin singlet in subsystem A. The possibility to get spin up or down particle are the same $1/2$ (1/C). So the number of configurations which obeys $(1,5)_{C=2}$ GPP is $\frac{20\times 19}{2^{2}}=95$. Up to now, one can claim that $\nu=1/5$ MATBCB under chiral limit hosts $C=2$ FCI phase.
@@
On the other hand, it is important to check if such $C=2$ FCI is robust when it deviates from the ideal condition [91]. For example, the influence of chiral ratio $\kappa=w_{AA}/w_{AB}$ [63, 62, 64]. When $\kappa\neq 0$, the lowest single particle bands will be broadened. The kinetic term will compete with projective interaction. And quasi-particle will generally feel a non-uniform Berry curvature. One can expect there is a critical chiral ratio where the FCI phase breaks down and phase transition happens. We will locate such phase transition according to PES gap [6]. In Fig. 8, we show the many-body energy spectrum and the PES of 10 fold degenerate mixed state under different chiral ratio $\kappa=0.5,0.625,0.6875,0.75$ while keeping other parameters unchanged. When $\kappa$ increases, there are 5 states approaching the original quasi-hole manifold at $\kappa=0$. As $\kappa$ reaches the critical value $\kappa_{c}$, these 5 states will collapse on lower quasi-hole manifold. Denoting $|\min(\xi(n=5,K^{A}=0\cdots 4))-\max(\xi(n=4,K^{A}=0\cdots 4),\xi(n=5,K^{A}=5\cdots 19))|$ as the PES gap, the PES gap closing phase transition can be interpreted as the violation of hard core condition. That means after phase transition, the particle with different pseudo spin can occupy the same orbital. So one has additional $\frac{20}{2^{2}}=5$ configurations. It is not difficult to understand this since when $\kappa$ increases, the kinetic term can be comparable with interaction at a critical value. By keeping track of the PES gap evolution about chiral ratio $\kappa$, we find that the PES gap decreases as $\kappa$ increases. And there is a PES gap closing at $\kappa_{C}\approx 0.7050$, which indicates a phase transition.
@@
To summarize this section, we have obtained the smoking gun evidence of the existence of $C=2$ FCI phase in $\nu=1/5,\kappa=0$ MATBCB via PES quasi-hole counting. The number of quasi-hole states under PES gap obeys $(1,5)_{C=2}$ GPP. As chiral ratio $\kappa$ increases, there will be PES gap closing at $\kappa_{c}\approx 0.7050$. The quasi-hole counting will change, which indicates the violation of the hard core condition. However, recall that in section  IV we get the spectrum flow with about $5-12{\rm meV}$ gap between ground state manifold and higher energy states. This gap is not large enough compared with the ground state manifold width $\approx 1.5{\rm meV}$. One may expect a weak external magnetic field or exchange interaction [92] is still needed to stabilize the FCI phase [13, 93], which will be left for future work.

### v2

@@ Abstract
We investigate a fractional Chern insulator (FCI) candidate arising from Moiré bands with higher Chern number $C=2$ on a magic angle twisted bilayer checkerboard lattice (MATBCB). There are two nearly flat low lying bands in the single particle energy spectrum under the first magic angle $\phi\approx 1.608^{\circ}$ and chiral limit. We find MATBCB hosts a nearly uniform Berry curvature distribution and exhibits tiny violation of quantum geometric trace condition in the first moiré Brillourin Zone (mBZ), indicating that there is a nearly ideal quantum geometry in MATBCB in single particle level. Turning on projected Coulomb interactions, we perform exact diagonalization and find a ten-fold ground state quasi-degeneracy in many body energy spectrum with filling fraction $\nu=1/5$. The ten-fold quasi-degenrate ground states further show spectra flow under flux pumping. By diagnosing the particle entanglement spectrum (PES) of the ground states, we obtain a clear PES gap and quasi-hole state counting consistent with Halperin spin singlet generalized Pauli principle, suggesting that a $C=2$ abelian fractional Chern insulator is realized in this system.
October 17, 2024
@@ I Introduction
These experiments have raised a great interest among theorists. Several theories have been proposed to interpret the FQAHE (zero magnetic field FCI) in $\rm{MoTe_{2}}$ [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and R5G [24, 25, 26, 27, 28, 29, 30]. Besides gapped phases, gapless phases beyond FQAHE, e.g., the composite fermi liquid phase, have also been found [31]. However, the experiments mentioned above all seem to focus on fractional filling $C=1$ Chern band which has well known lowest Landau level (LLL) correspondence. When the topological flat band host a higher Chern number, its potential FCI phase will not have a naive LLL correspondence [32, 33, 34, 35, 36]. What is more attracting, it is possible to realize non-abelian FCI phase [37, 32, 38, 39, 40, 41, 42, 20, 43, 44, 45, 46, 47, 48] in fractionally filled flat band with higher Chern number, which could be a building block for fault-tolerant quantum computation.
Motivated by above progress, it is valuable to search for higher Chern number FCI phase in real material or artificial system. In this manuscript, we theoretically study a $C=2$ abelian FCI candidate in magic angle twisted bilayer checkerboard (MATBCB) lattice model [49] with fractional filling $\nu=1/5$ of the lowest single particle conductive band. By using exact diagonalization (ED) in momentum space, we find there is a ten-fold ground state quasi-degeneracy in the many-body energy spectrum and the spectrum flow under adiabatic flux inserting. These ten-fold low energy states are the topological ground state candidates for putative $C=2$ FCI phase. From the perspective of many-body entanglement, we study the particle entanglement spectrum (PES) of these ten-fold ground states density matrix. We find that the quasi-hole counting under PES gap obeys the Halperin spin singlet generalized Pauli principle (GPP), which is a smoking gun evidence for the existence of $C=2$ FCI phase. This higher Chern number FCI phase may be realized in artificial quantum simulator such as Rydberg atom array [50], twisted optical lattice [51, 52, 53, 54], circuit QED system [55]. Up to now, we have noticed that several works have studied TBCB flat band [49, 56], or similar system like twisted square lattice [57, 58, 59, 60], twisted FeSe [61], strained moiré lattice with quadratic dispersion [62], and square lattice moiré heterostructure [63].
@@
Figure 2: (a) The Wilson loop winding of the lowest two flat bands near half filling. Since the single particle band is gapless at X, the Wilson loop is a $2\times 2$ matrix. The windings are corresponding to the two eigenvalues of the Wilson loop. One can read out the Chern number $C=\pm 2$. From the variation speed of the winding, one can believe that the Berry curvature distribution is uniform enough to host a FCI phase. (b) The Berry curvature given by Eq. (10). The k point density in the first mBZ is $74\times 74$. (c) The trace of quantum metric in the first mBZ using the same k point density. It almost has the same distribution with Berry curvature, which implies the single particle flat band Bloch function is close to a holomorphic function about $k=k_{x}+ik_{y}$ and ideal trace condition violation. The range of momentum $k_{x},k_{y}$ are the same $[0,2\pi]$ for (b) and (c).
@@ II Single particle moiré band structure of MATBCB
In this section we solve the Moiré Hamiltonian (free fermion hopping in momentum space) of MATBCB and show the flat band structure, as well as the non-trivial Chern number. To clarify, in the following we will use lower case $k$ to label the single particle momentum while the upper case $K$ denotes the many body momentum or total momentum. The two dimensional single particle wave vector is defined as $\textbf{k}=(k_{x},k_{y})$. The Greek symbol $\kappa=w_{AA}/w_{AB}$ represents chiral ratio to be introduced later. The Moiré Hamiltonian for a given momentum $H(k_{x},k_{y})$ expressed in reciprocal lattice basis is shown in Eq. (1):
@@
Up to now we have a preliminary picture of MATBCB in single particle level including single particle band and Wilson loop winding. In the next section, we will use quantum geometry indicators to inspect if MATBCB is favor to host FCI phase more carefully.
@@ III Quantum geometry indicators for fractional Chern insulator
In fact, there is still some subtlety when applying quantum geometry indicator to higher Chern number band, for which a recent generalization can be found in Ref. [69, 43]. Here for simplicity, out strategy is to transform the higher Chern number problem to familiar $C=1$ case. Fortunately, several recent works have shown that ideal higher Chern band can be decomposed into a set of $C=1$ ideal Chern bands [70, 71] under certain condition. Specially, when the Berry curvature of $C>1$ Chern band satisfies $F(\textbf{k})=F(\textbf{k}+\textbf{Q})$ with Q being some translation symmetry broken vector, the set of decomposed $C=1$ ideal Chern bands can form orthogonal basis. That means these $C=1$ bands can be independently filled. So the $C=1$ quantum geometry indicators can be used as usual. In our MATBCB case, $\textbf{Q}=(G_{1}/2,G_{2}/2)$. One can see this point later in detail. A more general FCI indicator may be the so- called vortexability [72, 73], which we leave for future work.
@@
For FCI phase, we mainly concern the Berry curvature $F$ and trace of quantum metric $Tr(g)$. In practice, one can use the following Wilson loop formalism Eq. (10) to get $F$, $Tr(g)$ simultaneously and accurately [43]. For the mBZ (torus $T^{2}$), there is no need to fix the gauge [74]. Caompared with simply using Eq. (8), Eq. (10) can better capture the non-linear character in $F$ and $Tr(g)$ especially when the k point is near the branch cut of Berry phase or the boundary of mBZ. Here we assume $dk_{x}=dk_{y}$.
$\begin{split}&[U_{r}(\textbf{k})]_{ab}=\langle u_{a}(\textbf{k})|u_{b}(\textbf{k}+d\textbf{k}_{r})\rangle,r=x,y.\\ &W(\textbf{k})=[U_{y}(\textbf{k})]^{-1}[U_{x}(\textbf{k}+dk_{y})]^{-1}U_{y}(\textbf{k}+dk_{x})U_{x}(\textbf{k})\\ &=\exp(-g_{k_{x}k_{x}}dk_{x}^{2}-g_{k_{y}k_{y}}dk_{y}^{2})\exp(iFdk_{x}dk_{y}),\\ &W=USV^{\dagger},\quad\exp(iFdk_{x}dk_{y})=UV^{\dagger},\quad S=\exp(-Tr(g)dk_{x}^{2}),\\ &C=\frac{1}{2\pi}\int_{mBZ}d^{2}kTr(F)=\frac{1}{2\pi i}\int_{mBZ}Tr(\ln(W(\textbf{k})))\\ &=\frac{1}{2\pi i}\int_{mBZ}\ln(\det(W(\textbf{k}))),\\ &F=\frac{\textbf{Im}(\ln(W))}{dk_{x}dk_{y}},\quad Tr(g)=-\frac{\textbf{Re}(\ln(W))}{dk_{x}^{2}}.\end{split}$ (10)
From Fig. 2, one can see that Berry curvature of $C=+2$ band (Wilson eigenvalue) and the trace of quantum metric are consistent, which means the ideal trace condition violation condition is well satisfied. The profile of Berry curvature is similar to the result given by Ref. [75], which indicates topological heavy fermion is also a potential competing phase. More importantly, one can see that the Berry curvature host the symmetry $F(\textbf{k})=F(\textbf{k}+\textbf{Q})$, $\textbf{Q}=(G_{1}/2,G_{2}/2)$. That means the $C=2$ band can be orthogonally decomposed into two $C=1$ bands. Therefore the original quantum geometry indicator for $C=1$ case is still valid in our $C=2$ MATBCB system. For MATBCB, we summarize the quantum geometry indicators as follow:
@@
    The standard deviation of Berry curvature is $\sigma[\eta]=6.0579$.
@@
    The trace condition violation is $T[\eta]=1.6357\times 10^{-7}$.
Comapared with MATBG [13, 68], MATBCB has more ideal bandwidth and trace condition violation but slightly worse Berry curvature deviation. We note that some recent works have proposed the concept of higher Berry curvature [76, 77, 78], which may develop to many-body version quantum geometry. Also the trace condition deviation can be generalized to local momentum k version [79]. How to apply them to moiré system will be an interesting problem.
Figure 3: The variation of trace condition violation $T[\eta]$ over chiral ratio $\kappa$ under $\Delta=1meV$ case. One can read out that $T[\eta]$ is monotonically increasing with $\kappa$, which indicates that non-zero chiral ratio can induce the instability of putative FCI phase.
@@ IV Many-body energy spectrum and spectrum flow for FCI candidate
Figure 5: (a) The many body energy spectrum for $\nu=1/5$ MATBCB under chiral limit. The discretization of the first mBZ is $4\times 5$. The number of mBZ is $7\times 7$ which is the size of quadratic nodes array. We will use the same parameters in the following simulation unless specified explaination. (b) The zoomed in of low energy window in (a). One can identify the ground state manifold with GSD=10 at total momentum sector $K=5,6\cdots 9$ and $K=15,16\cdots 19$. They are approximately related by vector $G_{1}/2$. The gap between ground state manifold and the higher excitation states is about 12 meV.
@@
Figure 6: (a) The spectrum flow under adiabatic flux insertion $k_{y}\to k_{y}+\Phi/N_{y}$ under chiral limit. (b) The zoomed in of low energy window in (a). One can examine that the states related by $G_{1}/2$ will keep almost two-fold degeneracy during flux inserting. The ten-fold quasi-degenerate ground states braid with each other as flux varying, while do not mix with higher energy states. These 10 states can be considered as the FCI topological ground state candidates. The gap between ground state manifold and higher energy states is about $5-12{\rm meV}$, While the width of ground state manifold is approximately $1-2{\rm meV}$.
@@ V The particle entanglement spectrum for MATBCB and quasi-hole counting
Figure 7: The PES of ten-fold ground state density matrix for chiral limit $\kappa=w_{AA}/w_{AB}=0$. The PES is labelled by the total momentum of subsystem A $K^{A}$. The spectrum almost gather at $\xi<16$ regime, while other data with higher values can be considered as noise. There is an apparent PES gap at $\xi\approx 6.5$. One can extract the PES quasi-hole counting at each $K^{A}$ (total momentum of subsystem A) sector under the PES gap. For $K^{A}=0,1\cdots 4$, there are 4 states per sector. While for $K^{A}=5,6\cdots 19$, there are 5 states per sector. The total number of quasi-hole is 95 (for $4\times 5$ momentum space size). The counting obeys the $(1,5)_{C=2}$ GPP based on Halperin spin singlet ansatz as argued in main text.
@@
From the above analysis, now one can come back and inspect the PES of $\nu=1/5$ MATBCB. In Fig. 7, we show the PES of the density matrix consists of 10 fold topological ground state candidates. There is an apparent PES gap at about $\xi=6.5$ as shown by the blue dashed line. One can check that there are 4 quasi-hole states under PES gap in each $K^{A}=0,1\cdots 4$ sector while 5 low quasi-hole states in each $K^{A}=5,6\cdots 19$ sector. The total number of quasi-hole state is 95 for $N_{x}=4,N_{y}=5,N_{A}=2$. Such quasi-hole counting is coincident with $(1,5)_{C=2}$ GPP. One can interpret this result as follows. In our case, $N_{x}=4,N_{y}=5,N_{e}=4,N_{A}=2$. One can enumerate the Halperin spin singlet in subsystem A. The possibility to get spin up or down particle are the same $1/2$ (1/C). So the number of configurations which obeys $(1,5)_{C=2}$ GPP is $\frac{20\times 19}{2^{2}}=95$. While for CDW phase, the quasi-hole counting is significantly different, which is given by $n=dC(N_{e},N_{A})$ [102, 42]. $d$ is the number of CDW order. Up to now, one can claim that $\nu=1/5$ MATBCB under chiral limit hosts $C=2$ abelian FCI phase. MATBCB can naturally realize the $C=2$ FCI without stacking $C=1$ FCI. For the later one, it is hard to guarantee the position of the valley is invariant when stacking.
@@
On the other hand, it is important to check if such $C=2$ FCI is robust when it deviates from the ideal condition [103]. For example, the influence of chiral ratio $\kappa=w_{AA}/w_{AB}$ [67, 66, 68]. When $\kappa\neq 0$, the lowest single particle bands will be broadened. The kinetic term will compete with projective interaction. And quasi-particle will generally feel a non-uniform Berry curvature. One can expect there is a critical chiral ratio where the FCI phase breaks down and phase transition happens. We will locate such phase transition according to PES gap [6]. In Fig. 8, we show the many-body energy spectrum and the PES of 10 fold degenerate mixed state under different chiral ratio $\kappa=0.5,0.625,0.6875,0.75$ while keeping other parameters unchanged. When $\kappa$ increases, there are 5 states approaching the original quasi-hole manifold at $\kappa=0$. As $\kappa$ reaches the critical value $\kappa_{c}$, these 5 states will collapse on lower quasi-hole manifold. Denoting $|\min(\xi(n=5,K^{A}=0\cdots 4))-\max(\xi(n=4,K^{A}=0\cdots 4),\xi(n=5,K^{A}=5\cdots 19))|$ as the PES gap, the PES gap closing phase transition can be interpreted as the violation of hard core condition. That means after phase transition, the particle with different pseudo spin can occupy the same orbital. So one has additional $\frac{20}{2^{2}}=5$ configurations. It is not difficult to understand this since when $\kappa$ increases, the kinetic term can be comparable with interaction at a critical value. By keeping track of the PES gap evolution about chiral ratio $\kappa$, we find that the PES gap decreases as $\kappa$ increases, shown in Fig. 9. And there is a PES gap closing at $\kappa_{C}\approx 0.7050$, which indicates a phase transition.
@@
To summarize this section, we have obtained the smoking gun evidence of the existence of $C=2$ FCI phase in $\nu=1/5,\kappa=0$ MATBCB via PES quasi-hole counting. The number of quasi-hole states under PES gap obeys $(1,5)_{C=2}$ GPP. As chiral ratio $\kappa$ increases, there will be PES gap closing at $\kappa_{c}\approx 0.7050$. The quasi-hole counting will change, which indicates the violation of the hard core condition. However, recall that in section  IV we get the spectrum flow with about $5-12{\rm meV}$ gap between ground state manifold and higher energy states. This gap is not large enough compared with the ground state manifold width $\approx 1.5{\rm meV}$. One may expect a weak external magnetic field or exchange interaction [104] is still needed to stabilize the FCI phase [13, 105, 106, 94], which will be left for future work.